# GALLa: Graph Aligned Large Language Models for Improved Source Code Understanding

### v2

@@ Abstract
GALLa: Graph Aligned Large Language Models
for Improved Source Code Understanding
Ziyin Zhang12thanks: Work done during Ziyin’s internship at Ant Group. The first two authors contributed equally.  Hang Yu2*  Shijie Li2  Peng Di2  Jianguo Li2thanks: Corresponding authors.  Rui Wang1† 1Shanghai Jiao Tong University 2Ant Group {daenerystargaryen, wangrui12}@sjtu.edu.cn, {hyu.hugo, lijg.zero}@antgroup.com
@@ 3.1.2 Adapter
Alternative to the cross-attention layer, the adapter may also be a multi-layer perception (MLP), as used by LLaVA (Liu et al., 2023), which applies projection independently to each node. The main difference between cross-attention and MLP is that cross-attention allows for information exchange between the nodes, while MLP is applied independently to each node. Thus, we choose cross-attention as the adapter in the main results, and experiment with MLP in the ablation studies in Section 4.4.
@@ 4 Experiments
In this section, we first discuss the details of our experiments (4.1, 4.2), and then provide the results in 4.3 and 4.4.
@@ 4.2.2 Training
For the first stage training of GNN and adapter, we train the model on the graph data for 15 epochs with learning rate 1e-4, 1K warmup steps, weight decay 0.1, 240 global batch size, AdamW optimizer (Loshchilov and Hutter, 2019), and ZeRO stage 2 (Rajbhandari et al., 2020). The training takes place on two machines, each equipped with 8 A100 80G GPUs.
@@ 4.3 Results
Training Trans (Ja2Py/Py2Ja) Clone Repair Sum (Java/Py/JS) Defect Avg MLP G2C 73.2/66.9 95.2 11.7 15.1/11.6/14.1 61.6 43.7 G2C + GQA 76.2/68.9 95.3 12.7 14.9/12.1/13.6 59.9 44.2 MagNet G2C 75.6/65.3 94.9 11.2 14.5/11.3/13.9 63.1 43.7 G2C + GQA 74.4/65.1 95.0 12.8 15.0/12.4/14.1 63.2 44.0
@@ 5 Conclusion
Limitations
As we pointed out in the related work section, all existing methods that attempt to combine structural graphs with code language models focus on small, BERT-style models. Since our work is the first to utilize structural graphs to enhance decoder-only LLMs in understanding source code, we are unable to find any existing methods for fair comparison, and thus can only fine-tune the LLMs directly on source code as a baseline. However, our work can serve as a baseline for any future research in graph-enhanced (or more generally, structure-aware) code LLMs.
Due to our limited computational budget, we mostly focused on LLMs around 1B-10B parameters (especially in the ablation studies) and only verified the effectiveness of GALLa on one 14B model. We leave the exploration of even larger structure-aware code LLMs to future work. Also due to the budget constraint, we mainly focused on multi-task training in our main experiments and did not conduct individual hyper-parameter searches for the five downstream tasks. However, we maintained a fair comparison between the baseline and the proposed methods.
Ethical Statements
All the datasets and models used in this paper are publicly available resources. We use them only for research purpose.
Acknowledgments
This paper is supported by Ant Group and the General Program of National Natural Science Foundation of China (62176153).
@@ Appendix B GraphQA Prompts
- Tracing the edges leads to confirming {parent}, a {parent_type}, as the parent of {node_type} {node}."
@@ Appendix D Example Outputs with Single-Task Finetuning
if (txt.contains(" " )) {
return Arrays.asList(txt.split(" " ));
} else if (txt.contains("," )) {
@@
if " " in txt:
return txt.split(" ")
elif "," in txt:
return txt.split(",")

### v3

@@ 3.1.2 Adapter
Alternative to the cross-attention layer, the adapter may also be a multi-layer perception (MLP), as used by LLaVA (Liu et al., 2023), which applies projection independently to each node. The main difference between cross-attention and MLP is that cross-attention allows for information exchange between the nodes, while MLP is applied independently to each node. Thus, we choose cross-attention as the adapter in the main results, and experiment with MLP in the ablation studies in Section $4.$
@@ 4 Experiments
In this section, we first discuss the details of our experiments (4.1, 4.2), and then provide the results in 4.3 and $4.$
@@ 4.2.2 Training
For the first stage training of GNN and adapter, we train the model on the graph data for 15 epochs with learning rate 1e-4, 1K warmup steps, weight decay 0.1, 240 global batch size, AdamW optimizer (Loshchilov & Hutter, 2019), and ZeRO stage 2 (Rajbhandari et al., 2020). The training takes place on two machines, each equipped with 8 A100 80G GPUs.
@@ 4.3 Results
Training Trans (Ja2Py/Py2Ja) Clone Repair Sum (Java/Py/JS) Defect Avg Baseline 73.8/53.3 93.5 10.1 14.6/11.3/13.8 61.1 41.4 MLP G2C 73.2/66.9 95.2 11.7 15.1/11.6/14.1 61.6 43.7 G2C + GQA 76.2/68.9 95.3 12.7 14.9/12.1/13.6 59.9 44.2 MagNet G2C 75.6/65.3 94.9 11.2 14.5/11.3/13.9 63.1 43.7 G2C + GQA 74.4/65.1 95.0 12.8 15.0/12.4/14.1 63.2 44.0
@@ Appendix B GraphQA Prompts
- Tracing the edges leads to confirming {parent}, a {parent_type}, as the parent of {node_type} {node}.”
@@ Appendix D Example Outputs with Single-Task Finetuning
if (txt.contains(” )) {
return Arrays.asList(txt.split(” ));
} else if (txt.contains(”,” )) {
@@
if in txt:
return txt.split(” ”)
elif ”,” in txt:
return txt.split(”,”)